OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 5x2 + 8y2 = 40.

OpenStudy (anonymous):

@paki @phi

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

do you know the conics formula to find these?

OpenStudy (anonymous):

\[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\]

OpenStudy (anonymous):

\[5x^{2} + 8y^{2} = 40\] you want the right side =1, so divide both sides by 40. \[\frac{ 5x^{2} }{ 40 } + \frac{ 8y^{2} }{ 40 } = \frac{ 40 }{ 40 }\] \[\frac{ x^{2} }{ 8 } + \frac{ y^{2} }{ 5 } = 1\]

OpenStudy (anonymous):

Now, \[a^{2} = 8 , b^2= 5\] \[a= 2\sqrt{2} , b = \sqrt{5}\]

OpenStudy (anonymous):

from here on, you can use this equation sheet to find foci, vertices, and center. Look under the section Ellipse, on the left side. https://www20.csueastbay.edu/library/scaa/files/pdf/Conics.pdf

OpenStudy (anonymous):

foci= (c,0) , (-c,0) so we need to find c \[c^{2} = a^{2} -b^{2}\] \[c^{2} = 8-5\] \[c= \sqrt{3}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\(\color{salmon}{\large\heartsuit}\large\text{You're Most Welcome! }\color{salmon}\heartsuit\) \(~~~~~~~~~~~~~~~\color{green}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\) \(~~~~~~~~~~~~~~~~~~~~\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar} \)